Integrate f(x,y)=x+y over the region bounded by y=x^2-1, x=2 and x=0.
1
Expert's answer
2021-10-04T14:07:48-0400

The sketch of the region is as shown in the figure below:





The integration of the function is evaluated as,


"\iint f(x,y)dA=\int_{1}^{2}\int_{0}^{x^2-1}(x+y)dydx"


"=\int_{1}^{2}[xy+(y^2/2)]_{0}^{x^2-1}dx"


"=\int_{1}^{2}(x^3-x+\frac {1}{2}(x^4-2x^2+1))dx"


"=\frac{1}{2}\int_{1}^{2}(x^4+2x^3-2x^2-2x+1)dx"


"=\frac{1}{2}[\frac{x^5}{5}+\frac{2}{4}x^4-\frac{2}{3}x^3-x^2+x]_{1}^{2}"


"=\frac{211}{60}"


Therefore, the double integral is "\iint f(x,y)dA=\frac{211}{60}"